OpenStudy (anonymous):

I just want to check my answers. Consider the problem: “James scored 78 on his first test. What must the least score on the second test be if his average score on both tests is to be more than 80?” To solve this problem, what must x represent? The least score?

OpenStudy (2bornot2b):

Let x represent james' score in the second test, so \[\frac{78+x}{2}=80\implies x=160-78\]

OpenStudy (anonymous):

bet you can do this in your head average of 78 and 82 is 80 so he must get better than 82%

OpenStudy (anonymous):

avg = (78 +x)/2 > 80 (78 +x) > 160 x > 160 -78 x > 82